(1 point) Find a non-zero vector x perpendicular to the vectors ✓ : 2 and ū -4 -2 =
Answer #1

find perpendicator vector ² Cinen TE 2 and ù= 3 Festit -2 ² = 2 x U= -L 2 3 L1 -2 3 = -[t2j-Ak ū= -1 +37= 28 2 : î [(2x-2)-(-